If an account has an APR of 3%, compounded monthly, what is the APY? Give your answer as a percentage (but leave out the % symbol) and round to 2 decimal places.

Suppose you put $1200 in a savings account at an APR of 5% compounded quarterly. How much interest was earned after 4 years? Round to the nearest cent and leave off the $ symbol.
Answer #1

1). APR = 3% compounded monthly

APY = (1+APR/n)^n - 1 = (1+0.03/12)^12 - 1 = 3.04%

2). PV = $1200

APR = 5% compounded quarterly

value after 4 years = PV*(1+r/n)^(n*t) = 1200*(1+0.05/4)^(4*4) = $1463.87

So interest earned = 1463.87 - 1200 = $263.87